The following are types of laws of Boolean except_________
Commutative
Associative
Distributive
None of the above
1
Expert's answer
2016-02-17T00:00:57-0500
d) None of the above

Some basic Boolean laws:
• Commutative Law
(a) A + B = B + A
(b) A B = B A
• Associate Law
(a) (A + B) + C = A + (B + C)
(b) (A B) C = A (B C)
• Distributive Law
(a) A (B + C) = A B + A C